### v1

@@ 6.1 Rotation sensors
A more direct use of persistent currents for rotation measurements can also be found. Taking advantage of the large control in experiments, a differential measurement of rotation around the critical rotation velocity at which a persistent current is created can be used to measure changes on rotational frequencies [120]. This, however, can have some limitations due to the hysteresis phenomenon [360] occurring around this critical point as well as unwanted phase slippage of coherent or incoherent nature [110, 428]. Nonetheless, the use of attractive interactions can give an edge on these measurements, where the angular momentum is fractionalized and therefore a current is created at much smaller changes of rotation, offering possibilities for creating absolute measures of rotation [120, 124]. Another approach for persistent current-based sensors can have been investigated using the analogue of the superconducting interference devices (SQUIDs), known as atomic SQUID or AQUID [133] (see Fig. 31). This can be exploited to create rotation sensors with entanglement-enhanced precision [429]. The superposition of orbital angular momentum states can also be used in a different manner to devise a rotation sensor. In a superposition of two orbital angular momentum states, a density minima appears due to the interference between the two counter-propagating modes. By precisely controlling the imbalance between the two states, one can sense rotations through the change on the frequency at which the nodal point of the density moves [430].
Figure 31: (a) Schematic of a double junction atomtronic SQUID with various rotation rates $\Omega_{ext}$ with $f$ being the frequency for the current bias generation and $N_{1},N_{2}$ corresponding to the number of atoms in each half of the ring. Arrows represent the movement of the junctions. (b) Change in normalized population difference $z$ as a function of the atom number at a rotation frequency of 00 Hz and 3 Hz. The red line corresponds to the optimal fit whilst the blue line indicates the critical atom number and the point of transition. Figures adapted from [133]
@@ 6.3 Quantum transport
Similarly, cold atoms pose a natural platform to study the interaction of transport and persistent currents. In particular, cold atoms can study quantum transport in a highly controlled and versatile way, allowing to probe regimes of transport not possible in other platforms [453, 455, 456]. For example, transport through a quantum ring depending on flux and persistent currents can be studied in novel ways using cold atoms [115, 116, 457, 232, 431]. In such a setup, a quantum ring threaded by flux $\Omega$ has two or more leads are attached, where the quantum transport through the ring is facilitated by the leads. One lead, commonly dubbed source, has a higher chemical potential, i.e. more atoms than the ring and other leads. Thus, atoms will start flowing from the source into the ring. The other leads, dubbed drain, have a lower chemical potential and thus the atoms in the ring start flowing into the drain leads. Now, the atoms propagating through the ring interact with the flux, which induces a path-dependent phase shift which can affect the transported current. The interplay between flux and quantum transport depends on the interaction strength between the atoms, as well as whether the atoms are fermions or bosons. For interacting bosons, the Aharonov-Bohm effect is washed out and the flux has little control over the current [115, 116]. The transport dynamics depends strongly on the ring-lead coupling, with weakly coupled rings showing oscillatory dynamics between source and lead which are modulated by the flux [115], while for strong coupling the Aharonov-Bohm effect is quickly washed out. In contrast, for non-interacting fermionic systems, the Aharonov-Bohm effect is present and the flux strongly affects the current. Curiously, we find that the dynamics of non-interacting fermions and hard-core bosons shows very different behaviors in a ring system with leads, while without leads the dynamics is exactly identically as the both species can be mapped onto each other.
@@
A different dynamics is found when the transport constitutes of low-energy excitations. Here, the system is prepared in the ground state, and a density wave is induced by slightly quenching the source lead. The incoming density wave is then either reflected and transmitted at the ring, where the transmission and reflection strongly depends on the flux. For zero flux, the density wave is transmitted with an amplitude of $4/3$ and a negative Andreev reflection, while for half flux quantum the density wave is not transmitted, but instead one observes a positive reflection [431].
@@
Quantum transport through a flux ring has been recently realized in experiment [459]. Experimentally, the lattice is realized in momentum space with non-interacting cold atoms subject to particle loss. The lattice sites are coupled via two-photon Bragg transitions induced by laser fields of different frequencies, whereas a coupling term with variable flux for the ring is induced by a four-photon process. The study find that transport through the ring coupled to bosonic condensate, which act as reservoir for atoms, is strongly modified by the flux.
@@ 6.4 Persistent current-based qubits and quantum phase slips
Figure 33: (a) Experimental realization of a ring-lattice potential with an adjustable weak link (red arrow). (b) Experimental realization of a ring-lattice potential with an adjustable weak link (red arrow). Measured intensity distribution with an azimuthal lattice spacing of 28 $\mu$m and a ring radius of 88 $\mu$m. The size of the structure is scalable and a lower limit is imposed by the diffraction limit of the focusing optics (c) Energy spectrum for $N_{p}=4$ attractively interacting bosons where a bound state has been formed and a single-site barrier is present with $\lambda=0.01$. We observe emerging avoided crossing forming new possible qubit states at lower rotational frequencies. Right panel corresponds to the current oscillations of a superposition of currents forming a qubit-like state created by quenching from $\Omega_{i}/\Omega_{0}=0$ to $\Omega_{f}/\Omega_{0}=0.125$ while the bad one is created by quenching to $\Omega_{f}/\Omega_{0}=0.175$. Figure adapted from [122, 138, 124]
From a theoretical point of view, two-level systems have been investigated both in continuum [137, 140] and lattice systems [138]. Important elements have been discussed such as the scaling of the gap, or possible implementations of gates. The high controllability also brings a large parameter space to explore for instance, the amplitude of persistent currents in presence of a barrier [121] (being one of the main mechanisms to couple different angular momentum states) or studies of AQUID with attractive interactions [124] (see Fig. 33(c)-(d)). Important issues on dissipation have also been explored in rings with weak impurities [109, 110]. Other proposals, inspired by the analogous circuit found in electronics (fluxonioums), include three weak links [142] where the links facilitating the function of the system as a qubit and enlarging the parameter space of the two-level effective dynamics.
Qubit-like states consisting of persistent currents strongly rely on the control one can have on those states. Coherent control of current states with present experimental techniques remains a challenge, as they typically involve using barriers or defects to control or imprint specific winding numbers. The change from on persistent current to another is known as “phase slips” [46]. Many investigations on incoherent phase slips exists, which account for the main process for which the current dissipates. Coherently controlling and understanding this process can allow for the engineering of states [114].
@@ 7 Remarks & Outlook
Persistent current can form the basis of quantum devices exploiting propagating matter-wave. Our first thought is towards compact quantum sensing [483, 150, 147, 425, 484, 485, 151, 152, 153, 486], that is clearly relevant in many different contexts including quantum labs in microgravity [487, 488, 489] and inertial navigation [490, 491, 492, 493, 494]. In this context, based on the lesson thought by high precision photonic interferometry [495, 496], an interesting direction to pursue would be exploring protocols employing cold atoms in non-classical states [497, 419, 286, 420, 146].
A second direction to explore would be studying integrated atomtronics and hybrid circuits. In the simplest schemes, such circuits can be made of suitable coupled toroidal degenerate gas characterized by given winding numbers. The ability to transfer winding numbers within such structures [114] encloses a great potential both for basic quantum science, for example providing experimentally feasible platforms for observing coherent quantum phase slips or macroscopic quantum coherence, and to conceive new concepts of quantum devices as atomic counterparts of Rapid Single Flux Quantum Logic superconducting circuits [498]. Ultracold atoms confined within optical cavities hybrid circuits [499] can open the way for non-destructive probing of ultracold matter through persistent current, which can be relevant both for basic investigations in quantum science [500, 501] and introduce novel avenues for quantum technology [502, 503]. However, challenges persist in achieving high-performance light-matter interactions [504, 505]and in the low density regimes, where quantum gas microscopy could be required [62, 506, 229].
@@ Acknowledgements
We thank A.J. Leggett, A. Osterloh, V. Singh and the Quantum Physics division at Technology Innovation Institute (Abu Dhabi) for many useful discussions. We acknowledge V. Ahufinger, H. Perrin, P. Pieri, A. Yakimenko for insightful comments to the manuscript.

### v3

@@ 6.1 Rotation sensors
A more direct use of persistent currents for rotation measurements can also be found. Taking advantage of the large control in experiments, a differential measurement of rotation around the critical rotation velocity at which a persistent current is created can be used to measure changes on rotational frequencies [126]. This, however, can have some limitations due to the hysteresis phenomenon [384] occurring around this critical point as well as unwanted phase slippage of coherent or incoherent nature [116, 455]. Nonetheless, the use of attractive interactions can give an edge on these measurements, where the angular momentum is fractionalized and therefore a current is created at much smaller changes of rotation, offering possibilities for creating absolute measures of rotation [126, 130]. Another approach for persistent current-based sensors can have been investigated using the analogue of the superconducting interference devices (SQUIDs), known as atomic SQUID or AQUID [139, 456] (see Fig. 31). This can be exploited to create rotation sensors with entanglement-enhanced precision [457]. The superposition of orbital angular momentum states can also be used in a different manner to devise a rotation sensor. In a superposition of two orbital angular momentum states, a density minima appears due to the interference between the two counter-propagating modes. By precisely controlling the imbalance between the two states, one can sense rotations through the change on the frequency at which the nodal point of the density moves [458]. Double ring configurations have also been considered, in particular, sided-coupled rings are sensitive to external rotations [120] and can be utilized to build rotation and acceleration sensors [459, 460]
Figure 31: (a) Schematic of a double junction atomtronic SQUID with various rotation rates $\Omega_{ext}$ with $f$ being the frequency for the current bias generation and $N_{1},N_{2}$ corresponding to the number of atoms in each half of the ring. Arrows represent the movement of the junctions. (b) Change in normalized population difference $z$ as a function of the atom number at a rotation frequency of 00 Hz and 3 Hz. The red line corresponds to the optimal fit whilst the blue line indicates the critical atom number and the point of transition. Figures adapted from [139].
@@ 6.3 Quantum transport
Similarly, cold atoms pose a natural platform to study the interaction of transport and persistent currents. In particular, cold atoms can study quantum transport in a highly controlled and versatile way, allowing to probe regimes of transport not possible in other platforms [483, 485, 486]. For example, transport through a quantum ring depending on flux and persistent currents can be studied in novel ways using cold atoms [121, 122, 487, 254, 461]. In such a setup, a quantum ring threaded by flux $\Omega$ has two or more leads are attached, where the quantum transport through the ring is facilitated by the leads. One lead, commonly dubbed source, has a higher chemical potential, i.e. more atoms than the ring and other leads. Thus, atoms will start flowing from the source into the ring. The other leads, dubbed drain, have a lower chemical potential and thus the atoms in the ring start flowing into the drain leads. Now, the atoms propagating through the ring interact with the flux, which induces a path-dependent phase shift which can affect the transported current.
The interplay between flux and quantum transport depends on the interaction strength between the atoms, as well as whether the atoms are fermions or bosons. For interacting bosons, the Aharonov-Bohm effect is washed out and the flux has little control over the current [121, 122]. The transport dynamics depends strongly on the ring-lead coupling, with weakly coupled rings showing oscillatory dynamics between source and lead which are modulated by the flux [121], while for strong coupling the Aharonov-Bohm effect is quickly washed out. In contrast, for non-interacting fermionic systems, the Aharonov-Bohm effect is present and the flux strongly affects the current. Curiously, we find that the dynamics of non-interacting fermions and hard-core bosons shows very different behaviors in a ring system with leads, while without leads the dynamics is exactly identically as the both species can be mapped onto each other.
@@
A different dynamics is found when the transport constitutes of low-energy excitations. Here, the system is prepared in the ground state, and a density wave is induced by slightly quenching the source lead. The incoming density wave is then either reflected or transmitted at the ring, where the transmission and reflection strongly depends on the flux. For zero flux, the density wave is transmitted with an amplitude of $4/3$ and a negative Andreev reflection, while for half flux quantum the density wave is not transmitted, but instead one observes a positive reflection [461].
@@
Recently, quantum transport through a flux ring has been realized in experiment [489]. Experimentally, the lattice is realized in momentum space with non-interacting cold atoms subject to particle loss. The lattice sites are coupled via two-photon Bragg transitions induced by laser fields of different frequencies, whereas a coupling term with variable flux for the ring is induced by a four-photon process. The study find that transport through the ring coupled to bosonic condensate, which act as reservoir for atoms, is strongly modified by the flux.
@@ 6.4 Persistent current-based qubits and quantum phase slips
Figure 33: (a) Sketch of the effective two-level system/qubit formed due to the energy splitting of the ground-state (black dotted lines) by a barrier. Insets corresponds to the qualitative form of the effective potential $V_{\mathrm{eff}}$ at effective flux $\Omega=0,\pi,2\pi$. (b) Experimental realization of a ring-lattice potential with an adjustable weak link (red arrow). Measured intensity distribution with an azimuthal lattice spacing of 28 $\mu$m and a ring radius of 88 $\mu$m. The size of the structure is scalable and a lower limit is imposed by the diffraction limit of the focusing optics. (c) Energy spectrum for $N_{p}=4$ attractively interacting bosons where a bound state has been formed and a single-site barrier is present with $\lambda=0.01$. We observe emerging avoided crossing forming new possible qubit states at lower rotational frequencies. Right panel corresponds to the current oscillations of a superposition of currents forming a qubit-like state created by quenching from $\Omega_{i}/\Omega_{0}=0$ to $\Omega_{f}/\Omega_{0}=0.125$ while the bad one is created by quenching to $\Omega_{f}/\Omega_{0}=0.175$. Figure adapted from [128, 144, 130].
From a theoretical point of view, two-level systems have been investigated both in continuum [143, 146] and lattice systems [144]. Important elements have been discussed such as the scaling of the gap, or possible implementations of gates. The high controllability also brings a large parameter space to explore for instance, the amplitude of persistent currents in presence of a barrier [127] (being one of the main mechanisms to couple different angular momentum states) or studies of AQUID with attractive interactions [130] (see Fig. 33(c)-(d)). Important issues on dissipation have also been explored in rings with weak impurities [115, 116]. Other proposals, inspired by the the analogous circuit found in electronics (fluxonioums), include three weak links [148] where the links facilitating the function of the system as a qubit and enlarging the parameter space of the two-level effective dynamics.
Qubit-like states consisting of persistent currents strongly rely on the control one can have on those states. Coherent control of current states with present experimental techniques remains a challenge, as they typically involve using barriers or defects to control or imprint specific winding numbers. The change from one persistent current to another is known as “phase slips” [50]. Many investigations on incoherent phase slips exists, which account for the main process for which the current dissipates. Coherently controlling and understanding this process can allow for the engineering of states [120].
@@ 7 Remarks & Outlook
Persistent current can form the basis of quantum devices exploiting propagating matter-wave. Our first thought is towards compact quantum sensing [513, 156, 153, 452, 514, 515, 157, 158, 159, 516], that is clearly relevant in many different contexts including quantum labs in microgravity [517, 518, 519] and inertial navigation [520, 521, 522, 523, 524]. In this context, based on the lesson taught by high precision photonic interferometry [525, 526], an interesting direction to pursue would be exploring protocols employing cold atoms in non-classical states [527, 446, 308, 447, 152]. A second direction to explore would be studying integrated atomtronics and hybrid circuits. In the simplest schemes, such circuits can be made of suitable coupled toroidal degenerate gas characterized by given winding numbers. The ability to transfer winding numbers within such structures [120] encloses a great potential both for basic quantum science, for example providing experimentally feasible platforms for observing coherent quantum phase slips or macroscopic quantum coherence, and to conceive new concepts of quantum devices as atomic counterparts of Rapid Single Flux Quantum Logic superconducting circuits [528]. Ultracold atoms confined within optical cavities hybrid circuits [529] can open the way for non-destructive probing of ultracold matter through persistent current, which can be relevant both for basic investigations in quantum science [530, 531] and introduce novel avenues for quantum technology [532, 533]. However, challenges persist in achieving high-performance light-matter interactions [534, 535]and in the low density regimes, where quantum gas microscopy could be required [66, 536, 537].
@@ Acknowledgements
We thank A.J. Leggett, A. Osterloh, V. Singh and the Quantum Physics division at Technology Innovation Institute (Abu Dhabi) for many useful discussions. We acknowledge V. Ahufinger, H. Perrin, P. Pieri, A. Yakimenko for insightful comments to the manuscript. LA acknowledges the Julian Schwinger Foundation grant JSF-18-12-0011.